1) Abe is thinking of buying a piece of commercial property as an investment. The property will cost $100,000. Abe believes that he can lease the property for $6,000 per year, payable at the beginning of every year, just as he buys it. He also believes that he can sell the property at the end of five years for $120,000. Calculate the present value (PV) of profits for Abe's business at each of the following discount rates: 8%, 9%, 10%, and 12%. Enter each of your calculated values in the table that follows and indicate whether or not Abe should buy the property at each of the discount rates. Finally, plot the PV for each of the discount rates rate on the graph and connect the plotted points with straight lines. At approximately what interest rate will Abe be indifferent between buying the property and not buying it?
1
Expert's answer
2019-09-20T09:27:01-0400

NPV(8%) = -100,000 + 6,000/1.08 + 6,000/1.08^2 + 6,000/1.08^3 + 6,000/1.08^4 + 6,000/1.08^5 + 120,000/1.08^6 = -423.38.

NPV(9%) = -100,000 + 6,000/1.09 + 6,000/1.09^2 + 6,000/1.09^3 + 6,000/1.09^4 + 6,000/1.09^5 + 120,000/1.09^6 = -5110.01.

NPV(10%) = -100,000 + 6,000/1.1 + 6,000/1.1^2 + 6,000/1.1^3 + 6,000/1.1^4 + 6,000/1.1^5 + 120,000/1.1^6 = -9518.41.

NPV(12%) = -100,000 + 6,000/1.12 + 6,000/1.12^2 + 6,000/1.12^3 + 6,000/1.12^4 + 6,000/1.12^5 + 120,000/1.12^6 = -17,575.61.

So, Abe should not buy the property at each of the discount rates.